OpenStudy (anonymous):

12(r - 3) = 6(r + 3)

OpenStudy (bahrom7893):

Okay so first divide everything by 6, what do u get?

OpenStudy (aroub):

12r-36=6r+18 12r-6r=36+18 6r=54 r=9

OpenStudy (bahrom7893):

Aww aroub don't just post the solution, get the asker a bit more involved

OpenStudy (aroub):

okay sorry but i showed the steps

OpenStudy (bahrom7893):

No still even if u post the solution. like instead of showing the first step, say: what will happen if u multiply out the parenthesis?

OpenStudy (bahrom7893):

I used to do that too, but polpak is really smart, he suggested this method.

OpenStudy (aroub):

okay!

OpenStudy (bahrom7893):

and once they respond say so now do this.. If they don't respond, move on.

OpenStudy (bahrom7893):

to the next question

OpenStudy (aroub):

yes it is , its very gd for learning

OpenStudy (saifoo.khan):

bahrom copied my method!

OpenStudy (bahrom7893):

saifoo go get lost.

OpenStudy (anonymous):

NO I LIKE UR STEPS AROUB

OpenStudy (bahrom7893):

yea because u don't have to do anything, but whatever. I'm not letting this snowball too

OpenStudy (aroub):

bahrom calm down!

OpenStudy (saifoo.khan):

wht the Heck? lol.

OpenStudy (anonymous):

WOW TRUE HADER ^ SHAIFF HALHA!!

OpenStudy (aroub):

haaha hadi! hadah zalameh 3anjad shayeh haloo!!

OpenStudy (anonymous):

HATER*

OpenStudy (anonymous):

OHHH LOL

OpenStudy (anonymous):

HOOWAH? ARABIE?

OpenStudy (aroub):

laa! 3ala ma azun hehe!

OpenStudy (anonymous):

HAHA!!

OpenStudy (aroub):

ur name is 7ana willa hana ??

OpenStudy (anonymous):

HANNAH

OpenStudy (aroub):

حنة??

OpenStudy (anonymous):

I CANT READ ARABIC LOL

OpenStudy (aroub):

oh! :(

OpenStudy (anonymous):

BUT YEA I THINK THATS HOW U SPELL MY NAME IN ARABIC

OpenStudy (aroub):

its actually a very nice name!

OpenStudy (anonymous):

thank you inti binit mish ahh?

OpenStudy (aroub):

yes u too right?

OpenStudy (anonymous):

yess lol

OpenStudy (aroub):

<3

OpenStudy (aroub):

sry i have to go to sleep now bye!:D <3